# The Basic Reproduction Number for Petri Net Models: A Next-Generation Matrix Approach

### v1

@@ 1 Introduction
In the context of epidemiology, the basic reproduction number ($R_{0}$) refers to the expected number of new cases that one infected individual is likely to cause in a population where the entire population is susceptible [1]. As such, $R_{0}$ is a fundamental concept in mathematical epidemiology for all communicable diseases when it comes to Ordinary Differential Equation (ODE) based models. Intuitively, $R_{0}$, represents the expected number of secondary infections caused by a single infected individual in a fully susceptible population [1, 2]. An accurate calculation of $R_{0}$ is therefore crucial to understanding an infectious disease’s potential spread and informing public health interventions[3]. Analyzing $R_{0}$ can give insights into the future dynamics of a particular disease based on the parameters, subpopulations, and constants that make up any given $R_{0}$ value. Researchers calculate $R_{0}$ in models to assess transmission dynamics, compare scenarios, and evaluate the effectiveness of control measures such as vaccination or social behavior [3]. Mathematical treatments of $R_{0}$ via the ”next generation method” (NGM) have been outlined by Diekmann [2, 4, 5] and Van den Driessche [6]. Although other approaches exist for calculating $R_{0}$, the next-generation method is one of the most commonly used for ODE models[3] and so we focus on it here.
@@
Applying the next generation method of calculating $R_{0}$ within the framework of Petri nets requires insights on how to bridge the gap between mechanistic model structure and formal spectral analysis. The contribution of this paper is introducing those insights in how the transition and transmission matrices are defined and characterizing $R_{0}$ in the context of Petri nets. Among the various formulations of the next generation method, the computational approach introduced by Diekmann et al. [2, 4, 5] is uniquely well-suited for implementation in Petri nets, as it explicitly constructs the next generation matrix from compartmental transitions, which naturally aligns with the token flow and transition structure of Petri nets. This compatibility allows for a direct mapping from Petri net dynamics to the linearized infection and transition matrices required for $R_{0}$ calculation.
@@ 3.1 Next-Generation Matrix for Ordinary Differential Equations
The method for calculating $R_{0}$ via the Next Generation Matrix (NGM) laid out by Diekmann et al. [2, 4, 5] defines $R_{0}$ as the spectral radius of the next generation matrix. Specifically, two matrices (transmission and transition matrices) are defined to find $R_{0}$, starting with $T$, which is defined as the transmission matrix. The transmission matrix outlines the production of new infections. Where elements within the matrices describe how people move from non-infected compartments to infected compartments. Next, they define $\Sigma$ as the transition matrix, which describes state changes due to individuals moving to other infected compartments, dying, or becoming immune. To then find $R_{0}$ we take the dominant eigenvalue of $(-T\Sigma^{-1})$ yielding $R_{0}=\varphi(-T\Sigma^{-1})$.
@@ 3.2 Next-Generation Matrix for Petri Nets
Since many of the papers we draw on for examples below use the van den Driessche and Watmough notation, we will lay out the Next Generation Matrix for Petri nets (NGMPN) following similar notation. First, think of places in Petri nets the same way that you might think about compartments in ODEs. Then we assume we have a Petri net model for a disease that fits the definition of VAPN or SPN laid out in Section 2. Assume that each transition can fire once per time step. Define the total amount infected places in the Petri net model as $T$. Thus, the infected places are defined as $I_{1},I_{2},...,I_{T}$. As a note, we define all places as $P$ and they are labeled $P_{1},...,P_{K}$, where $K$ is total number of places and $T\leq K$. We define $t_{N}$ as a transition with an input arc coming from a non-infected place and an output arc going to an infected place. With $w(t_{N_{j,...,k}},I_{i})$ being the sum of all arcs weights $j,...,k$ from arcs going from non-infected places going to an infected place $I_{i}$ for VAPN or the net sum of all rate equations $j,...,k$ for arcs from transitions with input arcs from non-infected places and output arcs going to infected places $I_{i}$ for SPN. Let $x=(x_{1},…,x_{n})^{t}$, with each $x_{i}\geq 0$, be the number of tokens in each place, being the Petri net equivalent of population in each compartment in an ODE. Then we define $\mathcal{F}_{i}(x)$ to be the rate at which tokens move into infected places. For the variable arc weight PN, this rate refers to the arc weights. This is represented in matrix form by
@@
We define $w(I_{i},t_{j,...,k})$ as the summation of arc weights of output arcs from infected place $I_{i}$ to transitions $t_{j,...,k}$ for VAPNs and the summation of rate equations of all arcs between infected place $I_{i}$ to transitions $t_{j,...,k}$. In other words, this represents adding up all arcs leaving infected place $I_{i}$. We define $w(t_{I_{i}},I_{j})$ as the arc weights of arcs from transition to infected places where the arcs leading into the transition come from infected place $I_{i}$, and this arc weight is equal to the arc weight of the arc from $t_{I_{i}}$ to $I_{j}$. Lastly, we define $w(t_{N_{m},...,N_{l}},I_{j})$ as the sum of the arc weights of arcs from transitions to infected place $I_{j}$ where the transitions $t_{N_{m},...,N_{l}}$ have no arcs leading into the transitions. We define $V^{-}_{i}$ to be the arc weight of tokens leaving infected place $i$ and $V^{+}_{i}$ to be the arc weight of tokens going into infected place $i$ outside of coming from non-infected places. In terms of what the terms $\mathcal{V}^{-}_{i}$ and $\mathcal{V}^{+}_{i}$ represent, it is the rate of people leaving the infected place $i$ and the rate of people going into infected place $i$ from anything other than coming from non-infected places, respectively. These terms combine to yield, $\mathcal{V}_{i}=\mathcal{V}^{-}_{i}-\mathcal{V}^{+}_{i}$. For Petri nets, this can be expressed through the matrix
$\mathcal{V}_{1,...,T}=\begin{bmatrix}w(I_{1},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{1})&-w(t_{I_{2}},I_{1})&\cdots&-w(t_{I_{T}},I_{1})\\ -w(t_{I_{1}},I_{2})&w(I_{2},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{2})&\cdots&-w(t_{I_{T}},I_{2})\\ \vdots&\vdots&\ddots&\vdots\\ -w(t_{I_{1}},I_{T})&-w(t_{I_{2}},I_{T})&\cdots&w(I_{T},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{T})\end{bmatrix}.$ (3)
@@
The disease free equilibrium (DFE) for a system with $K$ places is defined as ($P_{1}^{*},....,P_{K}^{*}$), the state at which the system remains absent of the disease. For the Petri nets the DFE can be found by first setting the token level of all infected places to zero. Then, the arc weights of arcs into and out of a given non-infected place $P_{i}$ are summed, with the the arc weights of arcs going out of $P_{i}$ being multiplied by -1. The resulting expression from these steps is then set to zero. From here, a system of equations is formed that can be solved for $P_{1}^{*},....,P_{K}^{*}$. Additionally, the total population, usually denoted as $N$, is defined as $\mathcal{N}+\varepsilon$, where $\mathcal{N}$ is the summation of tokens in all places, and $\varepsilon$ is the error from rounding in a discrete token level Petri net. In a continuous token level Petri net $\varepsilon=0$. From here, a system of equations is formed where the DFE token levels labeled of each non-infected place can be solved for.
@@
Then, using the Petri net based $F$ and $V$ values, the same Next Generation Matrix steps of finding the inverse of $V$ and finding the dominant eigenvalues of $FV^{-1}$ can be done to find the Petri net $R_{0}$. If the values of $F,V$ and the DFE of two systems are equal, then the resulting $R_{0}$ values are equal. However, the way that parameters interact within an ODE model and a PN model are different. We will show the process in detail with various examples with different functionality.
@@ 4 Case Studies
We have selected a range of ODE and corresponding PN example models to show how the Next Generation Matrix for Petri nets (NGMPN) approach to finding $R_{0}$. Each model either adds additional compartments/places or new dynamics to provide a range of examples for people to draw on if they wish to implement the NGMPN on a new model.
@@ 4.1 SIRS
The first step in finding the $R_{0}$ of the PN systems in Figure 1 is to obtain the Disease Free Equilibrium (DFE) by setting the token levels to zero in the infected places, setting the net arc weight of non-infected places to 0, and solving for the token values in non-infected places $(S^{*},I^{*},R^{*})=(S^{*},0,0)$. Then the transmission matrix starts with finding $\mathcal{F}_{i}(x)$, which can be found by looking at the arc weights of arcs going into the infected compartments as shown in Equation 2. Note that $i$ is with respect to the corresponding infected compartment, which we number. For this example, the $I$ compartment is 1.
@@
for both layouts. For the transition matrix, we start with $\mathcal{V}$, and we look at arc weights of arcs leaving the infected places as shown in Equation 3, yielding for both layouts
@@ 4.2 SEIR
The first step in finding the $R_{0}$ of the PN systems in Figure 2 is to obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the token values at the non-infected places $(S^{*},E^{*},I^{*},R^{*})=(\frac{\Pi}{\mu},0,0,0,0)$. Then the matrix $F_{i}(x)$ can be found by looking at the arc weights of arcs going into the infected compartments as shown in Equation 2. Note that $i$ is with respect to the corresponding infected compartment and the corresponding row in the matrix, which we number. For this example, the $E$ compartment is 1 and the $I$ compartment is 2.
@@
For $\mathcal{V}$, we look at the weights of arcs leaving the infected places and transitioning from one infected place to another. This again yields the same result for both Petri net options:
@@
This is analytically the same $R_{0}$ expression that can be found for the ODE using the next generation method.
@@ 5.2 Nonlinear System
As shown in Figure 9, the Petri net implementation of the nonlinear model produces $R_{0}$ values that closely match the algebraic solution, with RRMSE consistently below 0.1%. This high level of agreement demonstrates the accuracy and numerical stability of the Petri net-based next generation method, even in the presence of nonlinear transmission and recovery terms. These results support the validity of the approach for more complex and mechanistically detailed epidemiological models.

### v2

@@ 1 Introduction
In the context of epidemiology, the basic reproduction number ($R_{0}$) refers to the expected number of new cases that one infected individual is likely to cause in a population where the entire population is susceptible [1]. As such, $R_{0}$ is a fundamental concept in mathematical epidemiology for all communicable diseases when it comes to Ordinary Differential Equation (ODE) based models. Intuitively, $R_{0}$, represents the expected number of secondary infections caused by a single infected individual in a fully susceptible population [1, 2]. An accurate calculation of $R_{0}$ is therefore crucial to understanding an infectious disease’s potential spread and informing public health interventions[3]. Analyzing $R_{0}$ can give insights into the future dynamics of a particular disease based on the parameters, subpopulations, and constants that make up any given $R_{0}$ value. Researchers calculate $R_{0}$ in models to assess transmission dynamics, compare scenarios, and evaluate the effectiveness of control measures such as vaccination or social behavior [3]. Mathematical treatments of $R_{0}$ via the ”next-generation matrix method” (NGM) have been outlined by Diekmann [2, 4, 5] and Van den Driessche [6]. Although other approaches exist for calculating $R_{0}$, the next-generation method is one of the most commonly used for ODE models[3] and so we focus on it here.
@@
Applying the next-generation matrix method of calculating $R_{0}$ within the framework of Petri nets requires insights on how to bridge the gap between mechanistic model structure and formal spectral analysis. The contribution of this paper is introducing those insights in how the transition and transmission matrices are defined and characterizing $R_{0}$ in the context of Petri nets. Among the various formulations of the next-generation matrix method, the computational approach introduced by Diekmann et al. [2, 4, 5] is uniquely well-suited for implementation in Petri nets, as it explicitly constructs the next-generation matrix from compartmental transitions, which naturally aligns with the token flow and transition structure of Petri nets. This compatibility allows for a direct mapping from Petri net dynamics to the linearized infection and transition matrices required for $R_{0}$ calculation.
@@ 3.1 Next-Generation Matrix for Ordinary Differential Equations
The method for calculating $R_{0}$ via the Next-Generation Matrix (NGM) laid out by Diekmann et al. [2, 4, 5] defines $R_{0}$ as the spectral radius of the next-generation matrix. Specifically, two matrices (transmission and transition matrices) are defined to find $R_{0}$, starting with $T$, which is defined as the transmission matrix. The transmission matrix outlines the production of new infections. Where elements within the matrices describe how people move from non-infected compartments to infected compartments. Next, they define $\Sigma$ as the transition matrix, which describes state changes due to individuals moving to other infected compartments, dying, or becoming immune. To then find $R_{0}$ we take the dominant eigenvalue of $(-T\Sigma^{-1})$ yielding $R_{0}=\varphi(-T\Sigma^{-1})$.
@@ 3.2 Next-Generation Matrix for Petri Nets
Since many of the papers we draw on for examples below use the van den Driessche and Watmough notation, we will lay out the Next-Generation Matrix for Petri nets (NGMPN) following similar notation. To help in visualizing some of the concepts outlined in this process Figure 1 will be used as a basic VAPN model with non-infected places ($N_{1},N_{2}$) and infected places ($I_{1},I_{2},I_{3}$). Examples using SPNs will be given in the Case Studies section under various cases.
Figure 1: This VAPN is used as a guide to help visualize some of the concepts described in the Next-Generation Matrix for Petri Nets formulation. The places of this VAPN can be labeled $(P_{1},P_{2},P_{3},P_{4},P_{5})$ or equivalently $(N_{1},I_{1},I_{2},I_{3},N_{2})$, where the second label gives more description on the infectious state of the place.
First, think of places in Petri nets the same way that you might think about compartments in ODEs. Then we assume we have a Petri net model for a disease that fits the definition of VAPN or SPN laid out in Section 2. Assume that each transition can fire once per time step. Define the total amount infected places in the Petri net model as $T$. Thus, the infected places are defined as $I_{1},I_{2},...,I_{T}$. As a note, we define all places as $P$ and they are labeled $P_{1},...,P_{L}$, where $L$ is total number of places and $T\leq L$. We define $t_{N}$ as a transition with an input arc coming from a non-infected place and an output arc going to an infected place. With $w(t_{N_{j,...,k}},I_{i})$ being the sum of all arcs weights $j,...,k$ from arcs going from non-infected places going to an infected place $I_{i}$ for VAPN or the net sum of all rate equations $j,...,k$ for arcs from transitions with input arcs from non-infected places and output arcs going to infected places $I_{i}$ for SPN. Let $x=(x_{1},…,x_{n})^{t}$, with each $x_{i}\geq 0$, be the number of tokens in each place, being the Petri net equivalent of population in each compartment in an ODE. Then we define $\mathcal{F}_{i}(x)$ to be the rate at which tokens move into infected places. For the variable arc weight PN, this rate refers to the arc weights. This is represented in matrix form by
@@
Looking at Figure 1 we find that the elements utilized in $\mathcal{F}$ are outlined in red. This is continued in all the other Petri net models in this paper to make it easier to find what elements fit where by inspection. For $\mathcal{F}_{1}$ the element $w(t_{N_{j,...,k}},I_{1})=\beta_{1}$. For $\mathcal{F}_{2}$ the element $w(t_{N_{j,...,k}},I_{2})=\beta_{2}$. For $\mathcal{F}_{3}$ no transmission occurs directly into place $I_{3}$, so the element $w(t_{N_{j,...,k}},I_{3})=0$. Thus
$\mathcal{F}_{1,2,3}(x)$ $=$ $\begin{bmatrix}\beta_{1}&\beta_{1}&\beta_{1}\\ \beta_{2}&\beta_{2}&\beta_{2}\\ 0&0&0\end{bmatrix}.$
@@
We define $w(I_{i},t_{j,...,k})$ as the summation of arc weights of output arcs from infected place $I_{i}$ to transitions $t_{j,...,k}$ for VAPNs and the summation of rate equations of all arcs between infected place $I_{i}$ to transitions $t_{j,...,k}$. In other words, this represents adding up all arcs leaving infected place $I_{i}$. We define $w(t_{I_{i}},I_{j})$ as the arc weights of arcs from transition to infected places where the arcs leading into the transition come from infected place $I_{i}$, and this arc weight is equal to the arc weight of the arc from $t_{I_{i}}$ to $I_{j}$. Lastly, we define $w(t_{N_{m},...,N_{l}},I_{j})$ as the sum of the arc weights of arcs from transitions to infected place $I_{j}$ where the transitions $t_{N_{m},...,N_{l}}$ have no arcs leading into the transitions. We define $V^{-}_{i}(x)$ to be the arc weight of tokens leaving infected place $i$ and $V^{+}_{i}(x)$ to be the arc weight of tokens going into infected place $i$ outside of coming from non-infected places. In terms of what the terms $\mathcal{V}^{-}_{i}(x)$ and $\mathcal{V}^{+}_{i}(x)$ represent, it is the rate of people leaving the infected place $i$ and the rate of people going into infected place $i$ from anything other than coming from non-infected places, respectively. These terms combine to yield, $\mathcal{V}_{i}(x)=\mathcal{V}^{-}_{i}(x)-\mathcal{V}^{+}_{i}(x)$. For Petri nets, this can be expressed through the matrix
$\mathcal{V}_{1,...,T}(x)=\begin{bmatrix}w(I_{1},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{1})&-w(t_{I_{2}},I_{1})&\cdots&-w(t_{I_{T}},I_{1})\\ -w(t_{I_{1}},I_{2})&w(I_{2},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{2})&\cdots&-w(t_{I_{T}},I_{2})\\ \vdots&\vdots&\ddots&\vdots\\ -w(t_{I_{1}},I_{T})&-w(t_{I_{2}},I_{T})&\cdots&w(I_{T},t_{j,...,k})-w(t_{N_{m},...,N_{l}},I_{T})\end{bmatrix}.$ (3)
Looking at Figure 1 we find that the elements utilized in $\mathcal{V}$ are outlined in green. This is continued in all the other Petri net models in this paper to make it easier to find what elements fit in $\mathcal{V}$ by inspection. For $\mathcal{V}_{1}$ the elements $w(I_{1},t_{j,...,k})=\gamma_{1}$, $w(t_{N_{m},...,N_{l}},I_{1})=0$, $-w(t_{I_{2}},I_{1})=0$, and $-w(t_{I_{3}},I_{1})=0$. For $\mathcal{V}_{2}$ the elements $-w(t_{I_{1}},I_{2})=0$, $w(I_{2},t_{j,...,k})=\gamma_{2}$, $w(t_{N_{m},...,N_{l}},I_{2})=0$, and $--w(t_{I_{3}},I_{2})=0$. For $\mathcal{V}_{3}$ the elements $-w(t_{I_{1}},I_{T})=-\delta_{1}$, $-w(t_{I_{2}},I_{3})=-\delta_{2}$, $w(I_{3},t_{j,...,k})=\varepsilon$, and $w(t_{N_{m},...,N_{l}},I_{3})=0$. Thus,
$\mathcal{V}_{1,2,3}(x)$ $=$ $\begin{bmatrix}\gamma_{1}&0&0\\ 0&\gamma_{2}&0\\ -\delta_{1}&-\delta_{2}&\varepsilon\end{bmatrix}.$
@@
The disease free equilibrium (DFE) for a system with $L$ places is defined as ($P_{1}^{*},....,P_{L}^{*}$), the state at which the system remains absent of the disease. For the Petri nets the DFE can be found by first setting the token level of all infected places to zero. Then, the arc weights of arcs into and out of a given non-infected place $P_{i}$ are summed, with the the arc weights of arcs going out of $P_{i}$ being multiplied by -1. The resulting expression from these steps is then set to zero. From here, a system of equations is formed that can be solved for $P_{1}^{*},....,P_{L}^{*}$. Additionally, the total population, usually denoted as $N$, is defined as $\mathcal{N}+\varepsilon$, where $\mathcal{N}$ is the summation of tokens in all places, and $\varepsilon$ is the error from rounding in a discrete token level Petri net. In a continuous token level Petri net $\varepsilon=0$. The difference in total population between $N$ and $\mathcal{N}+\varepsilon$ comes into play when doing actual simulations of models not when finding $R_{0}$ using NGMPN method. From here, a system of equations is formed where the DFE token levels labeled for each non-infected place can be solved for.
@@
Then, using the Petri net based $F$ and $V$ values, the same Next-Generation Matrix steps of finding the inverse of $V$ and finding the dominant eigenvalues of $FV^{-1}$ can be done to find the Petri net $R_{0}$. If the values of $F,V$ and the DFE of two systems are equal, then the resulting $R_{0}$ values are equal. However, the way that parameters interact within an ODE model and a PN model are different. We will show the process in detail with various examples with different functionality.
@@ 4 Case Studies
We have selected a range of ODE and corresponding PN example models to show how the Next-Generation Matrix for Petri nets (NGMPN) approach to finding $R_{0}$. Each model either adds additional compartments/places or new dynamics to provide a range of examples for people to draw on if they wish to implement the NGMPN on a new model.
@@ 4.1 SIRS
The first step in finding the $R_{0}$ of the PN systems in Figure 2 is to obtain the Disease Free Equilibrium (DFE) by setting the token levels to zero in the infected places, setting the net arc weight of non-infected places to 0, and solving for the token values in non-infected places $(S^{*},I^{*},R^{*})=(S^{*},0,0)$. Then the transmission matrix starts with finding $\mathcal{F}_{i}(x)$, which can be found by looking at the arc weights of arcs going into the infected compartments as shown in Equation 2. These arcs are outlined in red in Figure 2. Note that $i$ is with respect to the corresponding infected compartment, which we number. For this example, the $I$ compartment is 1. From here we find $w(t_{N_{j,...,k}},I_{1})=\beta SI$ for the VAPN in Figure 2(a) and $w(t_{N_{j,...,k}},I_{1})=2\beta SI-\beta SI$ for the SPN in Figure 2(b). Thus,
@@
for both layouts. For the transition matrix, we start with $\mathcal{V}$, and we look at arc weights of arcs leaving the infected places as shown in Equation 3. These arcs are outlined in green in Figure 2. The terms $w(I_{1},t_{j,...,k})=\gamma I$ and $w(t_{N_{m},...,N_{l}}=0$ for both the VAPN in Figure 2(a) and the SPN in Figure 2(b). Thus yielding for both layouts
@@ 4.2 SEIR
The first step in finding the $R_{0}$ of the PN systems in Figure 3 is to obtain the DFE by setting the infected places token levels to zero, setting the net arc weight of non-infected places to 0, and solving for the token values at the non-infected places $(S^{*},E^{*},I^{*},R^{*})=(\frac{\Pi}{\mu},0,0,0,0)$. Then the matrix $F_{i}(x)$ can be found by looking at the arc weights of arcs going into the infected compartments, outlined in red in Figure 3 and as shown in Equation 2. Note that $i$ is with respect to the corresponding infected compartment and the corresponding row in the matrix, which we number. For this example, the $E$ compartment is 1 and the $I$ compartment is 2. The terms for the VAPN in Figure 3(a) within $\mathcal{F}_{1}$ are found by inspection to be $w(t_{N_{j,...,k}},I_{1})=\beta SI$ and within $\mathcal{F}_{2}$ to be $w(t_{N_{j,...,k}},I_{1})=0$. Similarly the terms for the SPN in Figure 3(b) within $\mathcal{F}_{1}$ are found by inspection to be $w(t_{N_{j,...,k}},I_{1})=\beta SI$ and within $\mathcal{F}_{2}$ to be $w(t_{N_{j,...,k}},I_{1})=\beta SI-\beta SI=0$. Thus,
@@
For $\mathcal{V}$, we look at the weights of arcs leaving the infected places and transitioning from one infected place to another, as outlined in green in Figure 3. Then for $\mathcal{V}_{1}$ the terms $w(I_{1},t_{j,...,k})=\eta E+\mu E$, $w(t_{N_{m},...,N_{l}}=0,I_{1})$, and $-w(t_{I_{2}},I_{1})=0$. For $\mathcal{V}_{2}$ the terms $-w(t_{I_{1}},I_{2})=-\eta E$, $w(I_{2},t_{j,...,k})=\alpha I+\mu I$, and $w(t_{N_{m},...,N_{l}},I_{2})=0$. Both $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are the same for the VAPN and SPN in Figures 3(a) and 3(b) respectively. Thus,
@@
This is analytically the same $R_{0}$ expression that can be derived for the ODE using the next-generation matrix method.
@@ 5.2 Nonlinear System
As shown in Figure 10, the Petri net implementation of the nonlinear model produces $R_{0}$ values that closely match the algebraic solution, with RRMSE consistently below 0.1%. This high level of agreement demonstrates the accuracy and numerical stability of the NGMPN method, even in the presence of nonlinear transmission and recovery terms. These results support the validity of the approach for more complex and mechanistically detailed epidemiological models.